In a classroom, there are 54 chairs. The number of chairs per row is three less than twice the number of rows. Find the number of rows and the number of chairs per row.
Answer by macston(5194)

.
c=chairs per row; r=number of rows
.
c=2r-3
.
(c)(r)=54
(2r-3)(r)=54
2r^2-3r=54
2r^2-3r-54=0
(2r+9)(r-6)=0
2r+9=0 OR r-6=0
r=-4.5 OR r=6
ANSWER 1: There are 6 rows (-4.5 makes no sense)
.
c=2r-3
c=2(6)-3=12-3=9
ANSWER 2: There are 9 chairs per row
.
CHECK:
(6 rows)(9 chairs/row)=54 chairs
54 chairs=54 chairs